your teacher makes a solution from LiOH. 5.03 grams of LiOH is mixed into 1.27 of water. From this information calculate the concentration of the solution
1
Expert's answer
2015-06-05T08:58:25-0400
The mass concentration of the substance: the mass of the substance divided by the volume of the mixture:
Cm = m/V (volume of LiOH is omitted in the total volume of solution is omitted. It is only 0.003 L comparing to 1.27 L of water)
Cm(LiOH) = 5.03/1.27 = 4 (g/L)
The molar concentration of the substance: amount of the substance divided by the volume of the mixture:
Cn = n/V = m/(M x V), where n=m(mass)/M(molar mass)
Cn = 5.03/(24 x 1.27) = 0.165 mol/L